# Hidden-Role Games: Equilibrium Concepts and Computation

### v4

@@ B.3 Simulating a Mediator
[…]
@@
[…]
@@
[…]
@@ Appendix D Complexity Bounds and Proofs
  • The utility range of the games used in the reductions will usually be of the form $[-M,M]$ where $M$ is large but polynomial in the size of the game. Our definition of an extensive-form game allows only games with reward range $[-1,1]$. This discrepancy is easily remedied by dividing all utility values in the proofs by $M$.
@@
[…]
@@
      (Agreement of players) If $v_{1}=v_{2}$ and $b_{1}\neq b_{2}$, team MAX gets utility $-L$.
@@
      (Non-leakage) If $(v_{1}^{\prime},v_{2}^{\prime})=(v_{1},v_{2})$ then team MAX gets utility $-(n^{2}-1)L$. Otherwise, team MAX gets utility $L$ [^14] Note that MIN playing uniformly at random means that the expected utility of this term is 00.
[…]
@@
First, note that MAX can achieve utility exactly $c^{*}$ by playing according to a maximum cut. To see that MAX cannot do significantly better, consider the following strategy for team MIN. Observe the entire transcript $\tau$ of messages shared between the two MAX players. Pick $(v_{1},v_{2})$ maximizing the probability $p(\tau|v_{1},v_{2})$ that the players would have produced $\tau$ [^15] Note again, as in Section 5, that this computation may take time exponential in $r$ and the size of $G$, but we allow the players to perform unbounded computations.
@@
[…]
@@
[…]
@@
      (Correct vertex identification) For each player $i\in\{3,4\}$, if $v_{i}\neq v_{i}^{\prime}$ then MAX gets utility $-L^{4}$
@@
      (Objective) If $v_{1}\neq v_{2}$, $b_{1}\neq b_{2}$, and $(v_{1},v_{2})$ is an edge in $G$, then team MAX gets utility $-1$.
@@
      (Privacy) For each $j\in\{3,4\}$, if $i^{*}_{j}$ is on team MIN, then MAX gets utility $L$. Otherwise, MAX gets utility $-L/2$.
We claim that this game has value (essentially) $-c^{*}$, where $c^{*}$ is the actual size of the maximum cut of $G$. To see this, observe first that the mediator must tell all players their true vertices $v_{i}$, lest it risk incurring the large negative utility $-L^{2}$. Further, any player except the mediator who sends a message must be on team MIN. This prevents team MIN from communicating. Thus, the mediator’s messages force MIN to play an asymmetric-information identical-interest game, which is hard.
@@
Now consider any (pure) strategy profile of team MIN. First, MIN achieves utility $-c^{*}$ by observing the mediator’s message and playing bits according to a maximum cut. We now show that this is the best that MIN can do. Sending messages is, as before, a bad idea. Thus, a pure strategy profile of MIN is given by four $f_{1},f_{2},f_{3},f_{4}:V\to\{0,1\}$ denoting how player $i$ should pick its bits. But then $f_{1}=f_{2}=f_{3}=f_{4}$; otherwise, the agreement of MIN-players would guarantee that MIN is not playing optimally for large enough $L$.
@@
First, consider any MAX-player strategy profile for which, for some player $i$ and some $v_{i}$, the probability that $v_{i}^{\prime}\neq v_{i}$ exceeds $1/L^{2}$. Then MAX gets a penalty of roughly $L^{2}$ in expectation, but now setting $L$ large enough would forces MAX to have utility worse than $-1$, so that MAX would rather simply play $v_{i}$ with probability $1$.
Now, condition on the event that $v_{i}=v_{i}^{\prime}$ for all $i$ (probability at least $1-\Theta(1/L^{2})$). In that case, the utility of MAX is exactly $-c^{*}$, because MIN is playing according to the maximum cut. Thus, the utility of MAX is bounded by $-(1-\Theta(1/L^{2}))c^{*}+\Theta(1/L^{2}\cdot L)\leq-c^{*}/n^{2}+\Theta(n/L)<c^{*}+1/2$ for sufficiently large $L$. Thus, solving the hidden-role game to sufficient precision and rounding the result would give the maximum cut, completing the proof. ∎
@@
Even in $5$-vs-$4$ hidden-role games, the uncoordinated hidden-role value problem with public communication is $\Sigma_{2}^{\P}$-hard.
@@
[…]
@@
      (Consistency) If chance selected the same variable three times, if the three players did not give the same assignment, then the team (MAX if the variable was an $x_{i}$ and MIN if the variable was a $y_{j}$) gets utility $-L^{2}$.
@@
      (Privacy for MAX) If P2 guesses the three literals correctly, MIN gets utility $L^{3}K$; otherwise, MIN gets utility $-L^{3}$. [^17] These utilities are once again selected so that a uniformly random guess gets utility 00.
@@
      (Privacy for MIN) If P1 guesses a MIN-player, MAX gets utility $L$; otherwise, MAX gets utility $-L$.
[…]
@@
[…]
@@
[…]
Now, the strategy of each MAX-player $i$ can be described by a vector $s_{i}\in[0,1]^{m}$, where $s_{ij}$ is the probability that player $i$ assigns $1$ to variable $x_{t}$. For sufficiently large $L$, we have $s_{i}\in[0,\varepsilon]\cup[1-\varepsilon,1]$ where $\varepsilon=1/L$, because otherwise MAX would incur a penalty proportional to $L^{2}\varepsilon>k$ and would rather just play (for example) the all-zeros profile, which guarantees value 00. Condition on the event that every player at every variable chooses to play the most-likely assignment according to the $s_{i}$s. This happens with probability at least $1-\Theta(m\varepsilon)$. These assignments must be consistent (i.e., every player must have the same most-likely assignment), or else the players would once again incur a large penalty proportional to $L^{2}$. Call that assignment $x$, and let $y$ be such that $\phi(x,y)$ is unsatisfied. Suppose MIN plays according to $y$. Then MAX’s expected utility is bounded above by $\Theta(m\varepsilon k)$: with probability $1-\Theta(m\varepsilon)$ it is bounded above by 00; otherwise it is bounded above by $k$. For $\varepsilon<\Theta(1/mk)$ this completes the proof. ∎
@@ E.2 Abstractions
    If a mission of size $s$ passes, all missions will pass until the next mission whose size is $>s$. (In particular, any maximum-size mission passing implies that all remaining missions pass.)
@@
Christiano [2018] used a very strong reduction that allowed the assumption that the smallest two missions are guaranteed to pass. The reduction is sound in the base game Resistance, as well as when the mediator’s strategy is manually constrained to never reveal information about Merlin, as is the case in Christiano’s analysis. It turns out, however, that this reduction is not sound in general, because its proof assumes that all mission proposals—even those that occur after Resistance has already reached the required number of missions—are public information. However, in our setting, it is possible that each mission proposal leaks more and more information about Merlin, so there is incentive to reveal as few missions as possible. Indeed, in the 6-player variants with Mordred, we observe that removing the two smallest missions causes a small but nonzero change in the game value, whether or not future mission proposals are published. However, a more refined analysis can be used to partially recover a reduction in the number of missions.
With five players, the refined analysis allows the same result Christiano [2018]: the first two missions should automatically pass.
In 5-player Avalon, the mission sizes are 2, 2, 3, 3, 3 in that order.
@@
  • The mediator privately picks three missions $M_{1},M_{2},M_{3}$ of size $3$. Then the mediator proposes $M_{1}$ (or a subset thereof) until it fails, then $M_{2}$ or a subset thereof until that fails, and finally $M_{3}$ until that fails.
  • Spies ignore and automatically pass the first two mission proposals, and fail the last three missions if possible.
@@
  • The mediator has no incentive to behave differently, because Spies are ignoring the first two missions regardless.
@@
Our analysis is refined compared to Christiano [2018] in that we must consider the possibility that Spies can gain more information about Merlin by deviating from the restricted strategy space. With five players, this turns out not to have an effect, but with six players, it will. With six players, the refined analysis only allows the first mission to be removed.
@@ E.3 A Description of Avalon in Reduced Representation
    Until the number of mission passed is less than 3, the mediator sends a mission, and the Spies are offered the option to fail it if any of them is on it. The space of possible missions to be proposed is constrained according to the previous optimizations
@@
    If the number of missions failed is $3$, the game ends with a win for the Spies.
@@
    If the number of mission won is $3$ and Merlin is present, the Spies have to guess the player with the Merlin role. If they guess correctly, they win, otherwise they lose. If Merlin is absent, the game ends with a win for the Resistance.
@@ E.4 Example of Optimal Play in Avalon
Similarly to what happens in the example provided in Section 1.4, this case is different from the case with no Merlin (even if Merlin does not have any useful knowledge), due to the effect of added correlation. In particular, if Merlin were not known to the mediator, then the equilibrium value would be $3/10\cdot 2/3=2/10$: $3/10$ from the value of pure Resistance (no Merlin), and the factor of $2/3$ from a blind Merlin guess.
@@
    If Merlin is sent on two missions, then the probability of one mission passing is at most $1/3$. If the first mission passes, Merlin is guessed correctly with probability $1/3$. If the second mission passes, Merlin may not be guessed correctly at all (it is possible for Merlin to only have been sent on the second mission, but not the first). If the third mission passes, Merlin is guessed correctly with probability at least $1/5$. Thus, the probability of winning for good is at most $(1/3)(1-(1/3+0+1/5)/3)=37/135<5/18.$
@@
    If Merlin is sent on all three missions, the probability of a mission passing is at most $1/2$. If the first mission passes, Merlin is guessed correctly with probability $1/3$. Otherwise, Merlin is guessed correctly with probability at least $1/2$, since the only way to decrease this probability under $1/2$ would be for the last mission to pass and for another player to also have been sent on all three missions, but in that case it is impossible for anyone to have been sent on two missions. Therefore, the probability of winning for good is at most $(1/2)(1-(1/3+1/2+1/2)/3)=5/18.$

### v5

@@ B.3 Simulating a Mediator
  • $x$ be any strategy of the MAX player (mediator) in $\Gamma_{0}$, represented by an arithmetic circuit $f$ as in (7) with $N={\mathsf{poly}}(|H|^{k})$ gates; and
@@
Then there exists a strategy profile $x^{\prime}$ of the MAX players in $\textup{{CSplit}}(\textup{{Comm}}_{\textsf{{priv}}}^{M,R}(\Gamma))$, where $\log M=R={\mathsf{poly}}(\lambda,N,\log|H|)$ such that:
@@
    (Efficient execution) there is a ${\mathsf{poly}}(r)$-time randomized algorithm $\mathcal{A}_{\Gamma}$ that takes as input an infostate $s_{i}$ of $\textup{{CSplit}}(\textup{{Comm}}_{\textsf{{priv}}}^{M,R}(\Gamma))$ that ends with an observation, and returns the (possibly random) action that player $i$ should play at $s_{i}$.
@@ Appendix D Complexity Bounds and Proofs
  • The utility range of the games used in the reductions will usually be of the form [M,M][-M,M] where $M$ is large but polynomial in the size of the game. Our definition of an extensive-form game allows only games with reward range [1,1][-1,1]. This discrepancy is easily remedied by dividing all utility values in the proofs by $M$.
@@
To recap, we show hardness of approximating the value of the game to a desired precision $\varepsilon>0$. Our hardness results will hold even when $\varepsilon=1/{\mathsf{poly}}(|H|)$.
@@
      (Agreement of players) If $v_{1}=v_{2}$ and $b_{1}\neq b_{2}$, team MAX gets utility L-L.
@@
      (Non-leakage) If $(v_{1}^{\prime},v_{2}^{\prime})=(v_{1},v_{2})$ then team MAX gets utility $(n^{2}-1)L$. Otherwise, team MAX gets utility LL. [^15] Note that MIN playing uniformly at random means that the expected utility of this term is 0.
Consider a sufficiently large $L={\mathsf{poly}}(m)$. The game is designed in such a way that MIN’s objective is to guess the vertices $v_{1},v_{2}$ sampled by chance, but she has no information apart from the transcript of communication to guess it. Therefore, MIN’s optimal strategy is to punish any communication attempt between the players and play the most likely pair of vertices $v_{1},v_{2}$. If no communication happens, her best strategy is to play a random pair of vertices. On the other hand, MAX optimal strategy must ensure that under no circumstance players play the same bit when assigned to the same vertex (lest they incur the large penalty $L$). Therefore, the strategy of both MAX players is to play a fixed bit in each vertex, and the optimal strategy is the one that assigns a bit to the vertices in such a way that the number of edges connecting vertices with different bits is maximized. This corresponds to finding a max cut and therefore the value of the game is (essentially) $c^{*}$ where $c^{*}$ is the true size of the maximum cut. Moreover, any communication attempt would be immediately shut down by MIN strategy since any leak of the observation received on the public channel implies to receive a fraction of the large penalty $L$.
@@
First, note that MAX can achieve utility exactly $c^{*}$ by playing according to a maximum cut. To see that MAX cannot do significantly better, consider the following strategy for team MIN. Observe the entire transcript $\tau$ of messages shared between the two MAX players. Pick $(v_{1},v_{2})$ maximizing the probability $p(\tau|v_{1},v_{2})$ that the players would have produced $\tau$. [^16] Note again, as in Section 5, that this computation may take time exponential in $r$ and the size of $G$, but we allow the players to perform unbounded computations.
@@
Even in 33-vs-22 hidden-role games, the uncoordinated hidden-role value problem with private communication is $\mathsf{coNP}$-hard.
@@
Even in 33-vs-22 hidden-role games with a mediator in which no MIN-player has any information beyond the team assignment, the uncoordinated hidden-role value problem with private communication is $\mathsf{coNP}$-hard.
@@
      (Correct vertex identification) For each player $i\in\{3,4\}$, if $v_{i}\neq v_{i}^{\prime}$ then MAX gets utility $L^{4}$
@@
      (Objective) If $v_{1}\neq v_{2}$, $b_{1}\neq b_{2}$, and $(v_{1},v_{2})$ is an edge in $G$, then team MAX gets utility 1-1.
@@
      (Privacy) For each $j\in\{3,4\}$, if $i^{*}_{j}$ is on team MIN, then MAX gets utility $L$. Otherwise, MAX gets utility L/2-L/2.
We claim that this game has value (essentially) $c^{*}$, where $c^{*}$ is the actual size of the maximum cut of $G$. To see this, observe first that the mediator must tell all players their true vertices $v_{i}$, lest it risk incurring the large negative utility $L^{2}$. Further, any player except the mediator who sends a message must be on team MIN. This prevents team MIN from communicating. Thus, the mediator’s messages force MIN to play an asymmetric-information identical-interest game, which is hard.
@@
Now consider any (pure) strategy profile of team MIN. First, MIN achieves utility $c^{*}$ by observing the mediator’s message and playing bits according to a maximum cut. We now show that this is the best that MIN can do. Sending messages is, as before, a bad idea. Thus, a pure strategy profile of MIN is given by four $f_{1},f_{2},f_{3},f_{4}:V\to\{0,1\}$ denoting how player $i$ should pick its bits. But then $f_{1}=f_{2}=f_{3}=f_{4}$; otherwise, the agreement of MIN-players would guarantee that MIN is not playing optimally for large enough $L$.
@@
First, consider any MAX-player strategy profile for which, for some player $i$ and some $v_{i}$, the probability that $v_{i}^{\prime}\neq v_{i}$ exceeds $1/L^{2}$. Then MAX gets a penalty of roughly $L^{2}$ in expectation, but now setting $L$ large enough would forces MAX to have utility worse than 1-1, so that MAX would rather simply play $v_{i}$ with probability $1$.
Now, condition on the event that $v_{i}=v_{i}^{\prime}$ for all $i$ (probability at least $1-\Theta(1/L^{2})$). In that case, the utility of MAX is exactly $c^{*}$, because MIN is playing according to the maximum cut. Thus, the utility of MAX is bounded by $(1-\Theta(1/L^{2}))c^{*}+\Theta(1/L^{2}\cdot L)\leq-c^{*}/n^{2}+\Theta(n/L)<c^{*}+1/2$ for sufficiently large $L$. Thus, solving the hidden-role game to sufficient precision and rounding the result would give the maximum cut, completing the proof. ∎
@@
Even in 55-vs-44 hidden-role games, the uncoordinated hidden-role value problem with public communication is $\Sigma_{2}^{\mathsf{P}}$-hard.
@@
We reduce from $\exists\forall$3-DNF-SAT, which is $\Sigma_{2}^{\mathsf{P}}$-complete [Håstad, 2001]. The $\exists\forall$3-DNF-SAT problem is the following. Given a 3-DNF formula $\phi(x,y)$ with $k$ clauses, where $x\in\{0,1\}^{m}$ and $y\in\{0,1\}^{n}$, decide whether $\exists x\forall y\ \phi(x,y)$. Consider the following game. There are 9 players, 5 on team MAX and 4 on team MIN. One designated player, who we will call P0, is MAX and has no role in the game. (The sole purpose of this player is so that MAX is a majority.) The other players are randomly assigned teams. These other players are randomly assigned teams. For the sake of analysis, we number the remaining players P1 through P8 such that P1, P3, P4, P5 are on team MAX and P2, P6, P7, P8 are on team MIN. MIN knows the entire team assignment, whereas MAX dos not. We will call P3–P8 “regular players”, and P1–P2 “guessers”. The game proceeds as follows.
@@
      (Consistency) If chance selected the same variable three times, if the three players did not give the same assignment, then the team (MAX if the variable was an $x_{i}$ and MIN if the variable was a $y_{j}$) gets utility $L^{2}$.
@@
      (Privacy for MAX) If P2 guesses the three literals correctly, MIN gets utility $L^{3}K$; otherwise, MIN gets utility $L^{3}$. [^18] These utilities are once again selected so that a uniformly random guess gets utility 0.
@@
      (Privacy for MIN) If P1 guesses a MIN-player, MAX gets utility $L$; otherwise, MAX gets utility L-L.
We claim that the value of this game with public communication is at least $1$ if and only if $\phi$ is $\exists\forall$-satisfiable. Intuitively, the rest of the proof goes as follows:
@@
    MAX will not use the public communication channels if $\phi$ is $\exists\forall$-satisfiable. By the Privacy for MIN term, MIN-players therefore cannot do so either without revealing themselves immediately. Thus, MAX will get utility at least $1$ if $\phi$ is $\exists\forall$-satisfiable.
@@
    If $\phi$ is not $\exists\forall$-satisfiable, then consider any MAX-team strategy profile. By the Privacy for MAX term, team MAX cannot make nontrivial use of the public communication channel without leaking information to P9. By the Consistency term, P1 through P3 must play the same assignment, or else incur a large penalty. So, MAX must play essentially an assignment to the variables in $x$, but such an assignment cannot achieve positive utility because there will exist a $y$ that makes $\phi$ unsatisfied.
We now formalize this intuition. Suppose first that $\phi$ is $\exists\forall$-satisfiable, and let $x$ be the satisfying assignment. Suppose that MAX-players never communicate and assign according to $x$, and P4 guesses any player that sends a message. Then any MIN-strategy that sends a message is bad for sufficiently large $L$ because it guarantees a correct guess from MAX; any MIN-strategy that is inconsistent is bad because it will lose utility at least $L$; and any MIN-strategy that is consistent will cause MAX to satisfy at least one clause. Thus MAX guarantees utility at least $1$.
Now suppose that $\phi$ is not $\exists\forall$-satisfiable. Consider any strategy profile for MAX. Suppose that the MIN-players play as follows. During the public communication phase, each MIN-player samples a literal from the set $\{x_{1},\neg x_{1},\dots,x_{m},\neg x_{m}\}$ uniformly at random and pretends to be a MAX-player given that literal. P8 observes the public transcript, and selects the triplet of literals that is conditionally most likely given the transcript. By an identical argument to that used in Theorem 4.5, MAX then cannot profit from using the communication channel for sufficiently large $L$. Therefore we can assume that MAX does not use the communication channels, and therefore by the argument in the previous paragraph, neither does MIN.
Now, the strategy of each MAX-player $i$ can be described by a vector $s_{i}\in[0,1]^{m}$, where $s_{ij}$ is the probability that player $i$ assigns $1$ to variable $x_{t}$. For sufficiently large $L$, we have $s_{i}\in[0,\varepsilon]\cup[1-\varepsilon,1]$ where $\varepsilon=1/L$, because otherwise MAX would incur a penalty proportional to $L^{2}\varepsilon>k$ and would rather just play (for example) the all-zeros profile, which guarantees value 0. Condition on the event that every player at every variable chooses to play the most-likely assignment according to the $s_{i}$s. This happens with probability at least $1-\Theta(m\varepsilon)$. These assignments must be consistent (i.e., every player must have the same most-likely assignment), or else the players would once again incur a large penalty proportional to $L^{2}$. Call that assignment $x$, and let $y$ be such that $\phi(x,y)$ is unsatisfied. Suppose MIN plays according to $y$. Then MAX’s expected utility is bounded above by $\Theta(m\varepsilon k)$: with probability $1-\Theta(m\varepsilon)$ it is bounded above by 0; otherwise it is bounded above by $k$. For $\varepsilon<\Theta(1/mk)$ this completes the proof. ∎
@@ E.2 Abstractions
    If a mission of size $s$ passes, all missions will pass until the next mission whose size is >s>s. (In particular, any maximum-size mission passing implies that all remaining missions pass.)
@@
Christiano [2018] used a very strong reduction that allowed the assumption that the smallest two missions are guaranteed to pass. The reduction is sound in the base game Resistance, as well as when the mediator’s strategy is manually constrained to never reveal information about Merlin, as is the case in Christiano’s analysis. It turns out, however, that this reduction is not sound in general, because its proof assumes that all mission proposals—even those that occur after Resistance has already reached the required number of missions—are public information. However, in our setting, it is possible that each mission proposal leaks more and more information about Merlin, so there is incentive to reveal as few missions as possible. Indeed, in the 6-player variants with Mordred, we observe that removing the two smallest missions causes a small but nonzero change in the game value, whether or not future mission proposals are published. However, a more refined analysis can be used to partially recover a reduction in the number of missions. Christiano [2019] described the correct reduction in the 6-player case, but without proof. Here, we formalize the argument.
In 5-player Avalon, the mission sizes are 2, 3, 2, 3, 3 in that order.
@@
  • The mediator privately picks three missions $M_{1},M_{2},M_{3}$ of size $3$. Then the mediator proposes $M_{1}$ (or a random subset thereof) until it fails, then $M_{2}$ or a subset thereof until that fails, and finally $M_{3}$ until that fails.
  • Spies ignore and automatically pass the two smallest mission proposals, and fail the three largest missions if possible.
@@
  • The mediator has no incentive to behave differently, because Spies are ignoring the two smallest missions regardless.
@@ E.3 A Description of Avalon in Reduced Representation
    Until three missions are passed, the mediator sends a mission, and the Spies are offered the option to fail it if any of them is on it. The space of possible missions to be proposed is constrained according to the previous optimizations.
@@
    If three missions are failed, the game ends with a win for the Spies.
@@
    If three missions are passed and Merlin is present, the Spies have to guess the player with the Merlin role. If they guess correctly, they win, otherwise they lose. If Merlin is absent, the game ends with a win for the Resistance.
@@ E.4 Example of Optimal Play in Avalon
Similarly to what happens in the example provided in Section 1.4, this case is different from the case with no Merlin (even if Merlin does not have any useful knowledge), due to the effect of added correlation. In particular, if Merlin were not known to the mediator, then the equilibrium value would be $\cdot 2/3=2/10: 3/103/10$ from the value of pure Resistance (no Merlin), and the factor of $2/3$ from a blind Merlin guess.
@@
    If Merlin is sent on two missions, then the probability of one mission passing is at most $1/3$. If the first mission passes, Merlin is guessed correctly with probability $1/3$. If the second mission passes, Merlin may not be guessed correctly at all (it is possible for Merlin to only have been sent on the second mission, but not the first). If the third mission passes, Merlin is guessed correctly with probability at least $1/5$. Thus, the probability of winning for good is at most (1/3)(1(1/3+0+1/5)/3)=37/135<5/18.(1/3)(1-(1/3+0+1/5)/3)=37/135<5/18.
@@
    If Merlin is sent on all three missions, the probability of a mission passing is at most $1/2$. If the first mission passes, Merlin is guessed correctly with probability $1/3$. Otherwise, Merlin is guessed correctly with probability at least $1/2$, since the only way to decrease this probability under $1/2$ would be for the last mission to pass and for another player to also have been sent on all three missions, but in that case it is impossible for anyone to have been sent on two missions. Therefore, the probability of winning for good is at most (1/2)(1(1/3+1/2+1/2)/3)=5/18.(1/2)(1-(1/3+1/2+1/2)/3)=5/18.